OpenStudy (anonymous):

A circle has its center at (-2, -3) and a radius of 3 units. What is the equation of the circle? (x + 2)2 + (y + 3)2= 9 (x + 2)2 + (y - 3)2= 9 (x - 2)2 + (y - 3)2 = 3 (x + 2)2 + (y + 3)2 = 3

OpenStudy (anonymous):

center radius form for the equation of a circle is: (x-h)^2 + (y-k)^2 = r^2 center is (h, k) an radius is r. identify your h, k, and r and put it in to the center radius form and simplify it.

OpenStudy (anonymous):

so would it be the second one?

OpenStudy (anonymous):

try again.

OpenStudy (anonymous):

well if the r^2 and r=3 then its the first one

OpenStudy (anonymous):

that's the one! :)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw